OpenStudy (anonymous):

My question got deleted :/ 4^-2x*4^x= 64 & 6^-2x*6^-x=1/216

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

let's hope there is no rude interruption this time!

OpenStudy (misty1212):

\[4^{-2x}\times 4^x=64\] first step is to add the exponents and get \[4^{-x}=64\]

OpenStudy (misty1212):

next write \(64\) as \(4^3\) and get \[\huge 4^{-x}=4^3\] you good from there?

OpenStudy (anonymous):

do I just dived 4 on both sides? then it turns to x=-3?

OpenStudy (misty1212):

lol no you don't "divide" but you do get \(x=-3\) it is because if the bases are the same, so are the exponents

OpenStudy (misty1212):

\[b^x=b^y\iff x=y\]

OpenStudy (misty1212):

so \[4^{-x}=4^3\iff -x=3\]

OpenStudy (misty1212):

next one is almost identical you want to try it?

OpenStudy (anonymous):

Oh, alright. thank you. and yes, for the next one do I turn 1/216 into 216^1?

OpenStudy (misty1212):

no write it as a power of 6

OpenStudy (anonymous):

So far I have 6^-x=1/216

OpenStudy (anonymous):

6^-x=6^3